OpenStudy (anonymous):

Steve can paint a room three times as fast as Billy. When they work together, Steve and Billy can paint a large room in 4 hours. How many hours would it take Billy to paint it by himself?

OpenStudy (anonymous):

@ash2326 @welshfella @mukushla

OpenStudy (anonymous):

a. 16/3 b. 3/16 c. 3 d. 16

OpenStudy (phi):

you need to use rate * time = 1 room rate is rooms/hour say billy works at rate "x" steve is 3 times faster, or 3x

OpenStudy (phi):

when they work together you can say the combined rate is x+3x and (x+3x)* 4 hours = 1 room 4x*4= 1 16x=1 x = 1/16 that is how fast billy works

OpenStudy (anonymous):

So 3/16 ?

OpenStudy (phi):

now use rate * time = 1 with rate = 1/16 to find the time

OpenStudy (phi):

they asked for How many hours would it take Billy to paint it by himself? (not the rate of steve... rate of steve is 3/16 (three times faster than bill))

OpenStudy (anonymous):

@phi So it woulf be 16/3